### v1

@@ 4 Experiments
Our experiments utilize the QuakeSet dataset [4] which contains 3,327 pairs of 2-channel Sentinel-1 imagery (VV and VH polarizations) of size $512\times 512$ sampled at different timesteps, $T_{1}$ and $T_{2}$, before and after an earthquake, respectively.
@@ 5 Conclusion
In this paper, we modified the problem of learning to regress earthquake magnitudes from satellite imagery as a ranking problem. Our results indicate that the use of margin ranking loss improves performance across all evaluated architectures to distinguish between areas affected by different magnitudes. Our method further benefits from requiring no additional data other than the existing magnitude labels, proving that the rankings between the predictions inside a batch can be employed to better add distance between samples with different magnitudes.

### v2

@@ 4 Experiments
Our experiments utilize the TorchGeo [18] implementation of the QuakeSet dataset [4] which contains 3,327 pairs of 2-channel Sentinel-1 imagery (VV and VH polarizations) of size $512\times 512$ sampled at different timesteps, $T_{1}$ and $T_{2}$, before and after an earthquake, respectively.
@@ 5 Conclusion
In this paper, we modified the task of learning to regress earthquake magnitudes from satellite imagery as a ranking problem. Our method benefits from requiring no additional data other than the existing magnitude labels, proving that the rankings between the predictions inside a batch can be employed to better add distance between samples with different magnitudes.